OpenStudy (anonymous):

du/dr=(1+sqrt(r))/(1+sqrt(u))_

OpenStudy (eyust707):

...

OpenStudy (across):

\[\frac{du}{dr}=\frac{1+\sqrt{r}}{1+\sqrt{u}}?\]

OpenStudy (anonymous):

yes

OpenStudy (eyust707):

that is a statment

OpenStudy (eyust707):

what would you like to do with that information?

OpenStudy (anonymous):

solve differential equation

OpenStudy (eyust707):

ahhhh

OpenStudy (eyust707):

its seperable..

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

this is how far i got but i want to know if its right. [u+u ^{3/2}=r+r ^{3/2}+C\]

OpenStudy (across):

That is supposed to be:\[u+\frac{2}{3}u^{3/2}=r+\frac{2}{3}r^{3/2}+c\]

OpenStudy (anonymous):

oh snap your right

OpenStudy (anonymous):

i forgot that step.-_- thanks:D

OpenStudy (anonymous):

but don't i have to get "u" by itself?

OpenStudy (across):

Usually, you should, but it is unfeasible to do that for this particular ODE (is inhumanely long).

OpenStudy (anonymous):

okay thankyou